Q6. A probability may be interpreted as a long-run _____________ frequency. (A) observational (B) relative (C)...

Q6.

A probability may be interpreted as a long-run _____________ frequency.

(A) observational

(B) relative

(C) experimental

(D) conditional

____________________________

Q7.

An important part of the customer service responsibilities of a cable company is the speed with which trouble in service can be repaired. Historically, the data show that the likelihood is 0.75 that troubles in a residential service can be repaired on the same day. For the first five troubles reported on a given day, what is the probability that all five will be repaired on the same day?

(A) .0010

(B) .6328

(C) .9990

(D) .2373

___________________________________

Q10.

Consider the experiment of tossing a fair coin three times and observing the number of heads that result (X = number of heads). What is the standard deviation for this distribution?

(A) 1.5

(B) 1.22

(C) 0.75

(D) 0.87
Answer #1

Answer :

Q (6).

The correct answer for the given statement is

option (B) is the correct

that is,

(B) relative

A probability may be interpreted as a long-run relative frequency

Q (7).

given data :-

the data show that the probability is 0.75

the first five troubles reported on a given day

now we need to find out the the probability that all five will be repaired on the same day?

= p( all five will be repaired on the same day )

= ( 0.75 )5

= 0.2373

Q (10).

given data :-

the experiment of tossing a fair coin = 3 times

n = 3

we can assume the probability of success = 0.5

p = 0.5

Now we need to find out the standard deviation

we know that

\sigma = sqrt( n * p * ( 1 - p ))

= sqrt ( 3 * 0.5 * ( 1 - 0.5 ))

= sqrt ( 3 * 0.5 * 0.5 )

=  sqrt ( 0.75 )

\sigma = 0.87